OpenStudy (socccerkid21):

What is the answer 3.86*6.47-3.97

OpenStudy (alivejeremy):

21.0042

Awolflover1 (awolflover1):

\[3.86\times6.47-3.97 \] \[3.86\times6.47=24.9742-3.97=? \] let's try not to get in trouble alive haha

OpenStudy (alivejeremy):

HA!